OpenStudy (anonymous):

Help?!?!Which of the following is true? A 21 is a composite number because one and twenty-one are its only factors. B 13 is a prime number because it is only divisible by one and itself. C 16 is a prime number because it is divisible by 4. D 28 is a composite number because 82 is a prime number. (Thanks<3)

OpenStudy (zzr0ck3r):

b... 21/3 = 7, 16/8 = 2 28/14 = 2

OpenStudy (anonymous):

Prime numbers can only be divided by itself and 1. Composite numbers are every other number that can be divided in ways other than 1 and itself. 13 is a prime number.

OpenStudy (zzr0ck3r):

1 is not prime also